OpenStudy (anonymous):

Identify the mathematical property that describes the following: -80 + 0 = -80

OpenStudy (btaylor):

Zero property of addition. x + 0 = x

OpenStudy (anonymous):

thsi is the additive property of zero, which says that if you add zero to something, literally nothing happen

OpenStudy (anonymous):

Additive property of zero. A number doesn't change when you add 0 to it. a+ 0 = a b+0 = b etc..